The concentrated sulfuric acid we use in the laboratory is 98.0 percent H2SO4 by mass. Calculate the molality and molarity of the acid solution. The density of the solution is 1.83g/mL


1
Expert's answer
2021-09-24T02:21:44-0400

Given:-

Mass of H2 SO4 = 98%


Means 98gm of H2 SO4 solute are present per 100gm of solution.


Hence mass of H2SO4 solute = 98gm.


Molar mass of H2SO4 solute =


molar mass = (1.008 g/mol × 2) + (32.07 g/mol × 1) + (16.00 g/mol × 4)


Molar mass = 98.086 g/mol


Mass of solution = volume x density

= 1000 x 1.83 = 1830 g


Mass of H2SO4 = 98.0/100 x 1830 = 1793.4 g


Moles of H2SO4 = mass/molar mass of H2SO4

= 1793.4/98.08

= 18.3 mol


Molarity = No. Of moles of H2SO4 solute / Volume of solution in l

=18.3/1

=18.3M


Therefore, the molarity of

solution H2SO4 is 18.3 M


Molality = moles of H2SO4/mass of water in kg


= 18.3/0.0366


= 500 mol/kg = 500m